OpenStudy (anonymous):

) Integrate ∫ (cos2x-cos2α)/ (cosx-cosα) dx, where α is independent of x.

OpenStudy (nikvist):

\[\int\frac{\cos{2x}-\cos{2\alpha}}{\cos{x}-\cos{\alpha}}dx=\int\frac{2\cos^2{x}-1-(2\cos^2{\alpha}-1)}{\cos{x}-\cos{\alpha}}dx=2\int\frac{\cos^2{x}-\cos^2{\alpha}}{\cos{x}-\cos{\alpha}}dx\]\[=2\int(\cos{x}+\cos\alpha)dx=2(\sin{x}+x\cos\alpha)+C\]

OpenStudy (anonymous):

Find the sum of the series 1² +3² +5² +-------to n terms.

OpenStudy (anonymous):

Find the sum of the series 1² +3² +5² +-------to n terms. who find my answer

OpenStudy (anonymous):

If the roots of the equation (a²+b²)x²-2(ac+bd)x+(c²+d²)=0 are equal, prove that a/b=c/d.